OpenStudy (anonymous):

find the derivative with respects to x: (4x^3)- (4ax^2)+ (a^2)x

OpenStudy (anonymous):

\[\frac{d}{dx}(4x^3-4ax^2+a^2x)\] I'm assuming a is a constant and not a variable in this case the derivative is: \[=12x^2-8ax+a^2\]

OpenStudy (anonymous):

12x^2-8ax+2a

OpenStudy (anonymous):

It wouldn't be 2a bc a is a constant so the derivative of (a^2)x= a^2

OpenStudy (anonymous):

u right nadeem, tenxz, IGNORE SHELLZ65 ANSWER